OpenStudy (richgall____):

Neil has been running a tutoring business since 2005. He charges a monthly fee for weekly tutoring sessions and a phone help line. Each year, he has increased his fee by the same amount. The table shows what Neil charged each customer for two given years of his business: Year Annual Tutoring Fee 2005 $1200 2008 $1350 A. What is the rate of change and initial value for Neil's business? How do you know? B. Write an equation in slope-intercept form to represent the fees that Neil charges each year.

OpenStudy (lolacole12):

ello

OpenStudy (richgall____):

hey

OpenStudy (lolacole12):

mate For the chart it's 1200 and 1350 for the prices and 2005 and 2008 for the years

OpenStudy (lolacole12):

Find the slope from the two given points: (0,1200) and (3,1350) \text{Slope } = \frac{ y_2 - y_1 } { x_2 - x_1 }= \frac{ 1350 - 1200}{ 3 - 0} = \frac{ 150}{ 3} =\boxed{\bf{ 50}} y = 50x + b Plug in the point: (0,1200) 1200 = 50(0) + b 1200 = b b = 1200 y = 50x + 1200 Part A. The rate of change is 50. The initial value is 1200. Part B. The equation in slope-intercept form is y = 50x + 1200.

OpenStudy (lolacole12):

the points 0 and 3, instead of 2005 and 2008, because 2005 was the first year he started his business. And therefore, 2008 was the 3rd year his business was running.

OpenStudy (richgall____):

wow that was fast

OpenStudy (richgall____):

Thankkssss.

OpenStudy (lolacole12):

u understand

OpenStudy (lolacole12):

ANY OTHER QUESTIONS MATEE

OpenStudy (richgall____):

thats it for now it was just this one

OpenStudy (lolacole12):

Here mite not see right but

OpenStudy (lolacole12):

:3